The sum of two numbers is 21 and their difference is 5. What are the number?
Found 2 solutions by rajgahir, mananth:
Answer by rajgahir(2)

Let the no. be x and y.

x + y = 21 ------eq(1)
x - y = 5 ------eq(2)

Add eq(1) and eq(2)
now we will eliminate y 'cause +y -y =0

2x = 26
x = 26/2
x = 13 now we have value of x , put the value of x in eq(2).
13 - y = 5
-y = 5 -13
-y = -8 i;e y = 8
So, the no.s are 13 and 8

Answer by mananth(16949)

x+y =21 x>y
x-y=5
2x=26
/2
x=13
plug the value of x in 1
13+y=21
y=21-8
we get y= 8